### v2

# Uncertainty in latent representations of variational autoencoders optimized for visual tasks

Abstract
Deep Generative Models (DGMs) can learn flexible latent variable representations of images while avoiding intractable computations, common in Bayesian inference. However, investigating the properties of inference in Variational Autoencoders (VAEs), a major class of DGMs, reveals severe problems in their uncertainty representations. Here we draw inspiration from classical computer vision to introduce an inductive bias into the VAE by incorporating a global explaining-away latent variable, which remedies defective inference in VAEs. Unlike standard VAEs, the Explaing-Away VAE (EA-VAE) provides uncertainty estimates that align with normative requirements across a wide spectrum of perceptual tasks, including image corruption, interpolation, and out-of-distribution detection. We find that restored inference capabilities are delivered by developing a motif in the inference network (the encoder) which is widespread in biological neural networks: divisive normalization. Our results establish EA-VAEs as reliable tools to perform inference under deep generative models with appropriate estimates of uncertainty.
Keywords Uncertainty  $\cdot$ Variational Autoencoders  $\cdot$ Probabilistic Inference  $\cdot$ Deep Generative Models  $\cdot$ Divisive Normalization
Introduction
Variational autoencoders (VAEs) have been established as immensely valuable tools to learn about the statistics of data in an unsupervised manner and to interpret it through inferring latent variables [1, 2, 3, 4]. Attractively, latent variables of a VAE, $\mathbf{z}$, can be considered as generative factors underlying data $\mathcal{P}(\mathbf{x})$, which directly links inferences in the model to the interpretation of observations. Intuitively, this generative process is summarized as a generative model $\mathcal{P}(\mathbf{x}|\mathbf{z})$, which is complete with a prior distribution over latent features, $\mathcal{P}(\mathbf{z})$. VAEs thus establish a highly principled approach to inference, approximating the true Bayesian posterior distributions $\mathcal{P}(\mathbf{z}|\mathbf{x})$ by a recognition model $\mathcal{Q}(\mathbf{z}|\mathbf{x})$. Notably, VAEs simultaneously estimate the most likely values of latent features and the uncertainty associated with these values. Indeed, an adequate representation of uncertainty is key in multiple scenarios, including decision making, and sensory integration [5], which rely on knowledge about the uncertainty of the sources [6]. Critically for the healthcare domain, where diagnostics involve integrating results from multiple tests and imaging modalities, an adequate estimate of uncertainty is required [7]. Consequently, capitalizing on the promise of VAEs requires reliable estimates of uncertainty.
Focusing on visual inference tasks, we identify four fundamental requirements towards principled inference using a salient variable of natural image statistics: contrast [8]. First, at zero contrast, which carries no information, the inferred posterior needs to be equal to the prior (Figure 1a). Second, with increasing contrast, observations become progressively more informative and the posteriors should become increasingly distinct from the prior. Accordingly, the average distance between the means of the posteriors and the prior, which we term signal mean, should increase (Figure 1a and b, top). Third, more informative observations yield increasing variance in posteriors means (Figure 1a and b, center), usually called signal variance. Finally, with increasing contrast, more informative observations denote decreasing uncertainty about the latent generative component, i.e. decreasing posterior variance, referred to as noise variance (Figure 1a and b, bottom). However, when training a VAE on natural images (Figure 1c), although contrast-dependence of the posterior mean matches the requirements, contrast-dependent changes in posterior variances are at odds with the promise of a faithful representation of uncertainty (Figure 1b).
Fig. 1: Characteristic properties of inferred posteriors as contrast is varied. a, Expected behaviour of the posteriors of a pair of latent variables in natural images when contrast is systematically increased. Three example images (green, purple and mustard frames) are presented at varying contrast levels. As contrast increases, the inferred posteriors for each image (colored solid lines) progressively deviate both from the prior (dashed line) and from one another. There are two different sources of variance in latent representations. The signal variance refers to how the posterior mean changes with the input. The noise variance (or posterior variance) refers to the remaining uncertainty after having observed a given stimulus. Similarly, we call signal mean to the average distance between the prior’s mean and the posteriors’ mean. The noise mean will be zero in our models, as is usually assumed for VAEs, without loss of generality. b, Cartoon of expected behaviour of signal mean, signal variance and noise variance as a function of contrast showing qualitative trends of these quantities. c,d, Signal mean, signal variance and noise variance of the inferred latent posteriors in natural image-trained VAE (c) and EA-VAE (d, respectively), as a function of image contrast (Methods). Prior mean and variance are shown in dotted black lines. e, Comparison of the inference and reconstruction process in the VAE (left) and EA-VAE (right). In the VAE a single pool of latent variables $\mathbf{z}$ is inferred, while in the EA-VAE there is an additional global latent variable $s$ which acts multiplicatively on the latents, $\mathbf{z}$. f, Example test image patches and their respective reconstruction through VAE and EA-VAE. g, Receptive filters of example latents in the VAE and EA-VAE. h, Inferred posterior mean of the scaling variable for individual patches (dots) in the EA-VAE model, as a function of the measured contrast of these images.
We argue that this signature of inaccurate inference is a consequence of the mismatch between the assumptions about the statistics of latent representations and those characteristic of generative factors underlying naturally occurring inputs. In its basic formulation, latent variables in VAEs are assumed to be either independent or linearly correlated. Natural stimuli, however, have been shown to display strong nonlinear dependencies between latents [9]. Thus, we propose a theoretically well-motivated and powerful update to the standard VAE architecture, the Explaining-Away-VAE (EA-VAE for short). EA-VAE incorporates a global multiplicative scaling variable $s$ in the generative model, which is inferred by the recognition model (Figure 1e) simultaneously with the traditional latent variables $\mathbf{z}$. This extension was inspired by the Gaussian Scale Mixtures (GSM) model [9], a simple yet powerful model of natural images for image compression and denoising, which has also been successful in interpreting behavioral and neural data in biological perception [10, 11, 12, 13]. The introduced scaling variable can replace correlated modulation of latent activations upon changes in image contrast, as recruiting the scaling variable can explain global changes in latent activations. Hence the term explaining-away in our EA-VAEs. In contrast to a standard VAE model, the EA-VAE produces contrast-dependent posteriors on natural images that match the fundamental requirements of probabilistic inference (Figure 1d).
In this work, we demonstrate that EA-VAEs address a broad set of issues concerning VAEs: the reliable representation of the posterior contributes to enhanced inference capabilities of EA-VAEs, and establishes parallels with computations in biological neural networks. First, we show that limited inference capabilities of VAEs are present in canonical computer vision domains such as handwritten characters recognition [14] and x-ray medical imaging [15], which are consistently remedied by EA-VAEs. Second, we demonstrate that the introduction of the scaling variable in the generative model shapes learned representations such that it promotes disentanglement of latent variables. Third, we demonstrate that EA-VAEs address challenges beyond contrast-dependent inferences, correcting inference for image manipulations that affect uncertainty and addressing a widely-known challenge of VAEs: the detection of out-of-distribution data. Fourth, we show the benefits of a more faithful uncertainty representation through information fusion. Finally, we identify biologically motivated properties of the recognition model of EA-VAEs that complement the inductive bias introduced in the generative model through the scaling latent variable. We demonstrate that improved inference in EA-VAEs is made possible by computations that bear the signatures of a widely known computational motif in biological neural networks: divisive normalization [16, 17]. The results open up multiple novel avenues of research, showing great promise for machine learning applications where uncertainty estimates are of importance, such as information fusion or anticipation of failure modes.
Results
Inference on natural images
To generate a first intuition for the role of the explaining-away variable $s$, we analysed natural image patches, where contrast inherently modulates uncertainty. In this setting, we train latent variable models to represent gray-scale square image patches (Figure 1f, top row and Methods). Sparse coding has historically been used to capture natural image statistics and recover classical simple-cell receptive filters and orientation selectivity as present in the primary visual cortex [18]. Motivated by these results, we investigate a version of VAEs, in which a similar representation has been shown to be induced through the use of Laplace (rather than Gaussian) priors [19, 20, 21]. We hence employ these models as a starting point for our analysis, though we will thoroughly analyze VAEs with Gaussian priors in later sections.
Latent representations
First, we train both a standard VAE and an EA-VAE under identical conditions: employing the same family of distributions, data and hyper-parameters. The only difference is that in the EA-VAE we select one of the latent variables to play the role of a scaling variable, multiplicatively affecting the output. The total number of latent variables is hence the same and the model complexity does not increase (Figure 1e). We find that both the VAE and the EA-VAE can be successfully trained on this task, showing a similar capacity to reconstruct natural images (Figure 1f and Figure S2a). Similarly, the models exhibited similar qualitative latent representations, with a fixed set of latents showing localized and oriented filters, and a smaller fraction showing unstructured filter properties (Figure 1g and Figure S2b). The fixed set of structured filters indicates that the system recruits as many units as needed given the effective dimensionality of the image space (Methods).
Latent uncertainty
Characterization of latents through filter properties informs us how the mean of the posterior is modulated by image properties. However, this analysis conveys little information about how the variance of the posterior, i.e. the uncertainty of inferences, is modulated by the input. In order to characterize the relationship between how stimulus content and contrast selectively contribute to latent activations, we characterize the posteriors through the signal mean, signal variance, and noise (or posterior) variance using natural image patches that display changes in their effective contrast level. These quantities are estimated in all cases as averages over the $\mathbf{z}$ latents (Methods). Both VAEs and EA-VAEs showed (as expected) increased signal mean and signal variance with higher contrast, as posteriors moved away from the prior and became more distinct (Figure 1, cf. b, c and d, top and middle panels). However, in standard VAEs, posterior uncertainty (noise variance) does not decrease with increasing contrast in the standard VAE (Figure 1c, bottom). Particularly, as contrast tends to zero, and local orientations become progressively harder to distinguish, posterior uncertainty fails to converge to the prior uncertainty. In contrast, EA-VAEs display a decreasing uncertainty for higher contrasts. Consistently with the assumptions of the generative model, as contrast decreases below the assumed pixel noise level, the reported noise variance in the EA-VAE saturates at the level of prior uncertainty (Figure 1d, bottom panel). This observation is also consistent with the fact that the latent $s$ is inferred as zero below this level (Figure 1h).
Improved inference of EA-VAE
Having established a first intuition for the role of explaining away variables in VAEs, we next explored whether the improved uncertainty representation in EA-VAEs may extend to other classical domains of computer vision. We hence turn to datasets widely used in the machine learning literature: MNIST [14] and ChestMNIST [15].
Uninformative observations
In their original form, MNIST and ChestMNIST feature images with standardized contrast, and consequently, contrast-dependence of the posterior is not relevant. However, standard computer vision tasks might benefit from an enhanced representation of uncertainty in settings other than low contrast situations.
Fig. 2: Characterization of posterior width for systematic image manipulations that affect inference uncertainty. a, Latent posterior width (mustard dots, with color matching that of the frame of average images) for uninformative images (top, mustard framed images) in the MNIST (left) and ChestMNIST (right) data sets. The distribution of noise posterior widths (std.) for in-distribution test images for standard VAE and EA-VAE are presented as violin plots. Prior uncertainty is shown as a reference (dashed line). b, Top left: Average posterior for images with the same label (’0’ in green and ’1’ in orange) in an MNIST-trained model. Top right: Gradual morphing between representative image samples from digit ’0’ and digit ’1’. Illustration of the expected changes of the posterior width is shown above. Bottom: Posterior widths for the morphed digits with standard VAE (left) and EA-VAE (right). c, Posterior width and the magnitude of the scaling variable of ChestMNIST-trained standard VAE and EA-VAE models for images increasingly corrupted either by Gaussian blurring (top) or additive pixel noise (bottom). d, Posterior width and the magnitude of the scaling variable for in-distribution (ID) and out-of-distribution (OOD) examples when trained on MNIST both for standard VAE (blue) and EA-VAE (red). Top: Examples of images from the same domain (gray frames) and different domains (mustard frames) as inputs, and reconstructions by the standard VAE (second row) and EA-VAE (third row) respectively. Bottom: uncertainty quantified by the posterior width and mean of scaling variable posterior for individual within-distribution or out-of-distribution images. As a reference, a dashed black line indicates the prior uncertainty.
First, to study the latent responses to an uninformative image, instead of a zero contrast image, we consider the ’average image’, obtained by a pixel-by-pixel average of all images from the dataset. As expected, this image resembles a blurry, unrecognizable digit (MNIST) or a slightly blurry x-ray (ChestMNIST) (Figure 2a). We characterize the latent uncertainty in the models measuring the noise standard deviation of the posterior and averaging along the $\mathbf{z}$ latent dimensions (Methods). In the case of MNIST, the VAE produced a latent uncertainty ($u_{{}_{\text{VAE}}}=0.17$) comparable to that of real handwritten digits ($u_{{}_{\text{ID}}}$) in the dataset [$p(u_{{}_{\text{ID}}}>u_{{}_{\text{VAE}}})=0.49$], and considerably below the unit standard deviation of the prior (Figure 2a left panel). Notably, the same measurement in the EA-VAE yielded $u_{{}_{\text{EA-VAE}}}=0.36$, This uncertainty is not only twice of that obtained for the VAE but also significantly higher than the uncertainty of actual samples from the distribution [$p(u_{{}_{\text{ID}}}>u_{{}_{\text{EA-VAE}}})=10^{-4}$]. A tendency similar to that observed on the MNIST dataset can also be seen on ChestMNIST-trained VAE and EA-VAE models [$p(u_{{}_{\text{ID}}}>u_{{}_{\text{VAE}}})=0.33$; $p(u_{{}_{\text{ID}}}>u_{{}_{\text{EA-VAE}}})=0.12$] (Figure 2a right panel). Note, that the average x-ray image resembles actual data samples more closely than the average digit image, which explains the less pronounced advantage of EA-VAE over the standard VAE. In summary, these results show that uncertainty estimates in the EA-VAE, but not in the standard VAE, distinguish uninformative images from actual images.
Morphing and uncertainty
When dealing with a domain containing distinct stimulus categories, the latent space is expected to adopt a non-overlapping representation in which inferences about individual images of a given category yield a low uncertainty estimate after learning. In contrast, inference is expected to reflect an increased uncertainty for images that might belong to multiple categories. Note, that as VAEs are trained in an unsupervised fashion, these do not rely on category labels during training. Still, we argue that distinct categories correspond to learned subdomains of the image space and morphing systematically drives images to domains of the pixel space with more limited exposure during training. Systematic changes in uncertainty due to ambiguous image categories can be investigated in the case of MNIST-trained models by assessing posterior uncertainty for images of simple digits and for images that are produced by merging two digits from different classes. We conducted this numerical experiment for each of the 45 pairs of different digits and measured the uncertainty along different levels of interpolations between the digits (Figure 2b, and Figure S4, for alternative pairings). We found a consistent modulation of uncertainty for the EA-VAE model, featuring a clear maximum at an intermediate interpolation level in all except one digit pairs (see also Methods). In contrast, no consistent peak was identified for the standard VAE model: only 31% of all cases featured a central maximum (14 out of 45 in total).
Image corruption and uncertainty
We next investigated the effect of noise-corruption on the representation of uncertainty. We relied on x-ray images from ChestMNIST, which we progressively corrupted, rendering them increasingly uninformative. We do so in two ways: by applying a Gaussian blur (Figure 2c top) and by adding Gaussian pixel noise (Figure 2c bottom) (Methods). In both cases the EA-VAE displays increasing uncertainty as the image quality deteriorates. This starkly contrasts with the behavior of standard VAEs, where latent uncertainty remains almost constant at low values even as the original image is no longer recognizable from the corrupted one (cf. blue and salmon lines, Figure 2c).
Out of distribution detection
Given that EA-VAEs present a robust improvement in uncertainty representation for uninformative, corrupted and interpolated images within each dataset, we investigated whether the introduced inductive bias of EA-VAEs could also contribute to enhanced out-of-distribution (OOD) detection. Specifically, we investigated whether uncertainty estimates of models reflect OOD stimulus presentations by reporting increased uncertainty. Using all the previously employed datasets and their corresponding VAE and EA-VAE trained models, we compare internal representations and reported uncertainties for within-distribution (ID) vs. OOD images (Figure 2d and Figure S5).
Using models trained on MNIST, we evaluated performance both ID (MNIST test partition) and OOD (pixel-shuffled MNIST, FashionMNIST, ChestMNIST, and natural images). Two key differences emerged between the VAE and the EA-VAE. First, reconstructed OOD images from the VAE were often interpreted as digits (Figure 2d top panels), while the EA-VAE produced characteristically blurry reconstructions. Second, the EA-VAE consistently reported higher values of latent uncertainty for OOD images than for images whose statistics matches that of the training distribution (Fig. 2d middle panel). Indeed, the EA-VAE reports uncertainty levels much closer to the prior when presented with OOD examples (further analysis in Methods and Supporting Information). These two facts are a clear indication that the generative model in standard VAEs reports unjustified certainty about its input. The EA-VAE, on the other hand can naturally deal with OOD examples, flagging with high uncertainty cases for which it was not trained on.
Characterization of latent representations
Next, we investigated the role of the explaining-away variable, $s$. We observed that the magnitude of $s$ is higher for MNIST images than for out of distribution examples (Figure 2d). Moreover, as expected given its multiplicative role, at natural image-trained EA-VAEs $s$ grows monotonically with image contrast (Figure 1h), which generalized across different choices of the distribution of $s$ (Figure S3). These results suggest that the explaining away variable $s$ could be playing the role of a signal-to-noise ratio estimate, with low values corresponding to an interpretation that the image is mostly noise, and consequently the posteriors for other latents relax towards the prior.
To corroborate these intuitions, we trained VAEs and EA-VAEs on extended versions of MNIST, ChestMNIST and FashionMNIST [22], where contrast was allowed to vary (Methods, referred to as cMNIST, cChestMNIST and cFashionMNIST, respectively). In both domains we found that the signal mean grows unbound with contrast in VAEs, while it quickly saturates for EA-VAEs (Figure 3a). Conversely, the explaining away variable was strongly modulated by contrast in the EA-VAE for both of these extended data sets (Figure 3b). This suggests that $s$ contributes to the disentanglement of stimulus content from contrast in EA-VAE.
Fig. 3: Characterization of learned representations for different training data sets. Beyond natural image patches, contrast-augmented versions of the MNIST and Fashion MNIST databases are shown. a, Latent posterior signal mean as a function of the measured contrast of a large set of test images for both the VAE and the EA-VAE models. b, Inferred posterior mean of the scaling variable for individual images (dots) in the EA-VAE model, as a function of the measured contrast of these images. Average mean shown as solid line. c, Direct comparison of the learned representation of an cMNIST-trained standard VAE (blue) and an EA-VAE (red) both featuring a three-dimensional latent space. Top panels: Cross-sections of the latent space (see labels above panels). Dots show the contrast of images generated from individual states of the latent space, with dot size proportional to the contrast of the generated image. Note that cross sections for the EA-VAE correspond to different levels of the scaling variable. Bottom panel: Contrast of images generated from the three-dimensional latent space of the standard VAE model at a fixed distance from the origin. The fixed distance was 2SD of the prior distribution.
We sought to understand how EA-VAE’s inductive bias shapes its learned representation. For this, we trained three-dimensional latent space versions of VAE and EA-VAE models, whose latent space can be directly visualized. Images generated from a grid in the latent space of VAE show strong modulations of image contrast in all latents (Figure 3c, top), with equal-contrast images residing approximately on a sphere (Figure 3c, bottom). In contrast, images generated from the regular latent variables of the EA-VAE ($z_{1}$ and $z_{2}$) showed homogenous contrast at any given value of the explaining away latent, $s$ (Figure 3c, middle). In summary, the explaining away variable contributes to the disentanglement of contrast from other latent factors in EA-VAEs, which does not naturally occur in standard VAEs (though see [23]).
From latent to predictive uncertainty
Next, we investigated whether classification based on the inferred latent features benefits from the learned uncertainty representation. Optimal information fusion relies on computing with proper uncertainty estimates, and VAEs are often employed for dimensionality reductions as a previous step to classification. Hence, when multiple sources need to be combined, such as in medical imaging, improved uncertainty estimates could be beneficial.
To investigate this, multilayer perceptrons (MLPs) were trained to classify MNIST digits using samples from the VAE (or EA-VAE) posterior latent representations (Methods). Two experiments evaluated the impact of these representations on information fusion and out-of-distribution examples. First, using a set of morphed digits, we compared the predictive uncertainty of MLPs, quantified as the entropy of the output distributions. The EA-VAE demonstrated significantly higher uncertainty (entropy) for intermediate images compared to the VAE ($p_{\text{value}}<10^{-15}$). (Figure 4a and Figure S6) Second, testing on out-of-distribution inputs (ChestMNIST and pixel-shuffled MNIST), the EA-VAE again yielded higher entropy predictions than the VAE (Fig. 4b).
These results show that differences in the properties of latent uncertainty representations between VAEs and EA-VAEs also translate to differences in output uncertainty, when those representations are used downstream for a supervised task.
Fig. 4: Characterization of a MLP trained for classification task of MNIST handwritten digits through latent posterior representations of VAE and EA-VAE. a, Behaviour of trained MLP when gradually morphing between image samples from digit ’2’ to digit ’5’. Top panels: Histograms of output predicted probabilities averaged for different samples of digits ’2’ and ’5’ when trained with standard VAE (left) and EAVAE (right) latent representations. Bottom panel: Entropy of predicted probability distribution as a function of combination weight of labels. b, Behaviour of trained MLP when testing with out of distribution chestMMNIST (Top) and pixel-shuffled MNIST (Bottom) images. Left columns: Histograms show the distribution of predicted probabilities of labels of an example image, averaged for different samples of the posterior distribution of that image. The entropy of that averaged distribution is computed for both models. Right column: Entropy of predicted probability distribution for MLP trained with standard VAE representation vs EA-VAE representations. Each yellow dot represents a single x-ray image. Black dot in mean value. Identity in dark grey slashed line. Shading corresponds to the upper limit determined by the uniform distribution.
Signatures of divisive normalization in the recognition model
So far, we have demonstrated that the inductive bias in the generative model contributes to more accurate inference. Next, we investigated the recognition model to understand the computations that enable it to accomplish improved inference.
Contrast-related changes induce a dependency in the joint statistics of linear filters, referred to as bow-tie dependency: pairwise conditional distributions display a characteristic bow-tie shape (Figure 5a). As VAEs assume independence of latent variables, we draw inspiration from neuroscience, where divisive normalization was proposed to reclaim independence [10] (see Methods).
Fig. 5: Divisive normalization implemented by the recognition model. a, Conditional distribution ($p(L_{1}\,|\,L_{2},x)$) of linear filter responses of a pair of example latent variables over natural images. Intensity of histogram is proportional to the probability. Linear filter responses are calculated as dot product between an image and the receptive field of the neuron (Methods). Vertical and horizontal axes correspond to $L_{1}$ and $L_{2}$ response intensities. Light and dark colored bars correspond to two central and two flanking quadrants of $L_{2}$ activation, respectively. b, Non-linear dependence of latent posteriors. Dependence is characterized by 100 randomly chosen pairs of latents using natural image posteriors. Standard deviation of the conditional distribution of posterior means are shown close to zero activation of the conditioned latent variable (two central quadrants, marked with matching color as panel a) and at high-intensity activation (two flanking quadrants, color as on a). c, Distribution of learned weights of the divisive normalization model for the standard VAE (blue) and EA-VAE (red) models. Dark line shows the average of five fits, individual fits are shown with light lines. d, Evaluation of divisive normalization through contrasting linear responses with posterior means in the standard VAE (blue) and EA-VAE (blue) models. Dots represent responses of individual latents for a particular natural image, colors distinguish responses to individual natural images, color lightness set according to image contrast. Lines show linear fits to the responses. Slope of the fit is designated as the normalization index, such that a smaller index corresponds to a decreased dependence of the posterior on the linear response, i.e. stronger divisive normalization. Normalization indices, and contrasts of the five example images are shown in the legend. e, Normalization index in the standard VAE (blue) and EA-VAE (blue) models as a function of the normalization factor of the divisive normalization model, i.e. the Euclidean mean posterior means. Dots correspond to individual images, light colors correspond to images with contrast levels below observation noise (Methods). f, Normalization index in the standard VAE model (blue) and EA-VAE model (red) as a function of image contrast. Dots represent the normalization index for a particular natural image. The vertical line signals the set observation noise.
We trained the divisive normalization model (Equation 9, Methods) on both the standard VAE and the EA-VAE. As the scaling variable explains contrast-related changes, we expected that the bow-tie dependency is reduced in EA-VAE. To analyze this, we computed the conditional distributions of the responses of latent pairs over natural images (Figure 5a) and calculated the width of the conditional distribution both at low-, and high-level activations of the conditioned latent (Figure 5b). High-level activations were characterized by higher variance in both the standard VAE and EA-VAE, but the difference was reduced in the EA-VAE (Figure 5b). We found that the reduced bow-tie dependence could be explained by stronger divisive normalization in EA-VAE, as reflected by stronger weights learned by the divisive normalization model (Kolmogorov-Smirnov test: $D=0.598\pm 0.153$, Figure 5c).
We also assessed divisive normalization by measuring the degree of suppression of individual latents depending on the activity of the remaining latent variables. Suppression was measured through the normalization index, which is based on the relationship between the linear responses and posterior means: a small normalization index indicated stronger suppression relative to the response expected from a linear model (Figure 5d). We compared the response intensity of the latents with the normalization index for both the EA-VAE and standard VAE models (Figure 5e). While the EA-VAE displayed consistently higher normalization with more intense latent activations, the standard VAE showed an opposite tendency, indicating that divisive normalization was characteristic of EA-VAE but not the standard VAE. Note that observation noise limits the quality of inference, as images with contrast lower than the level of observation noise are considered merely as noise by the generative model. Increased contrast results in higher signal mean (Figure 1c-d) and therefore contrast might have a direct effect on the strength of divisive normalization. Indeed, contrast-dependence of the normalization index identified a strong effect of divisive normalization in the EA-VAE model, but not in the standard VAE (Figure 5f).
In summary, we identified signatures of divisive normalization in EA-VAEs, supporting the conclusion that the inductive bias shapes the learning of the recognition model of the EA-VAE.
Discussion
Deep generative models are praised for establishing a coherent framework for probabilistic machine learning [24, 25], and VAEs are particularly appealing as they establish a normative tool for learning of (and inference in) generative models of arbitrary complexity [26]. However, our investigations consistently revealed systematic problems in VAEs uncertainty representations, which we argued stems from ignoring non-linear statistical dependencies characteristic of natural and artificial images. Motivated by generative models which employ multiplicative variables such as the GSM [9], we incorporated a global scaling latent variable to VAEs. In these models, which we call EA-VAEs, the simultaneous high activation of a large number of local variables in the VAE can be explained away by an increase in the global multiplicative variable. This allows EA-VAEs to effectively capture non-linear dependencies in the data and to reliably represent uncertainty in a wide range of visual tasks. Our analyses revealed the role of divisive normalization in the enhanced inference engine of EA-VAEs, providing further support for the benefit of neuroscience-inspired ideas in artificial intelligence [27].
We argue this work constitutes a strong example of how the interaction between ML and neuroscience can be mutually beneficial. Indeed, the use of deep learning methods as modeling tools in neuroscience has proven highly effective in recent years [28], and goal driven approaches in particular continue to push boundaries in our understanding of the links between perception and neural responses [29, 30, 31]. However, previous work has mostly focused on point estimates of most likely inferences. From a theoretical point of view, the Bayesian framework provides a natural means to incorporate the notion of uncertainty in the context of perception [32]. Since exact Bayesian inference is often intractable, two main approaches have consolidated over the years in order to perform approximate inference: variational inference and sampling-based inference. In this sense, uncertainty representation has been studied in the past in recurrent models of the cortex implementing sampling-based inference [13], to our knowledge, there is still an important gap when it comes to the study of the quality of uncertainty representations in variational models, such as VAEs. This is highly relevant given the tendency of deep CNNs to suffer from calibration problems, resulting in inaccurate uncertainty reports [33].
In this work, we present substantial evidence that posteriors inferred by VAEs systematically struggle to represent uncertainty in a meaningful way in a variety of settings, such as image corruption, interpolation, and out-of-distribution detection. In contrast, our results show that EA-VAEs systematically outperform standard VAEs when it comes to uncertainty representations. EA-VAE posteriors tend to more strongly revert to the prior when images become less informative by either corruption or interpolation, and report high uncertainty for out-of distribution examples. Moreover, we showed how benefits in latent uncertainty representations also translate to improved output uncertainty estimates basing their decisions in latent activations.
Moreover, as shown in Ref. [13], artificial neural systems optimized for full Bayesian inference beyond point estimates, for models with explaining away variables such as the GSM, result in solutions with mean responses which precisely recover divisive normalization, a canonical computational mechanism in the cortex [16]. It is hence possible that, as with divisive normalization in the brain, explaining away variables in artificial neural architectures may play an important role in multiple domains. Indeed, although originally motivated by the study of cortical representations for visual perception in the context of natural images, we have shown equally beneficial properties in standard computer vision datasets such as MNIST and ChestMNIST. While further work is required to assess these benefits in other non-visual modalities, we believe EA-VAEs constitute a promising model type for a wide range of scenarios. Inference models which can readily provide a meaningful representation of uncertainty are crucial in a plethora of fields, such as the medical domain, where optimal information fusion relies on weighting multiple information sources by their relative uncertainties [7].
Methods
Variational Autoencoders
To obtain approximate solutions to the process of Bayesian probabilistic inference, we trained VAEs to simultaneously learn the generative model through a decoder, and the inference model through an encoder. For any input $\mathbf{x}\in\mathbb{R}^{M}$, the encoder calculates the parameters of a variational approximation of the posterior, $p(\mathbf{z}|\mathbf{x})$, through $q_{\phi}(\mathbf{z}|\mathbf{x})$, where $\mathbf{z}\in\mathbb{R}^{D}$ is the vector of latent features with an assumed prior $p(\mathbf{z})$. After taking a sample $\mathbf{z}$ from $q_{\phi}(\mathbf{z}|\mathbf{x})$ the autoencoder attempts to reconstruct the input by synthesizing $\hat{\mathbf{x}}$. The reconstruction is produced via the decoder, which corresponds to the likelihood $p_{\varphi}(\mathbf{x}|\mathbf{z})$ of the generative model (Figure 1e, left). The likelihood can be considered as a mapping $f(\mathbf{z})$ from $\mathbf{z}$ to $\mathbf{x}$, which parametrizes the mean of the likelihood for a given $\mathbf{z}$. In our study, depending on the pixel distribution of the data, the likelihood was constructed in terms of a normal or a Bernoulli distribution, with independent pixels in both cases. In the case of a normal likelihood the variance of the distribution, the observation noise $\boldsymbol{\sigma}_{obs}^{2}$, was not learned.
The objective function to learn the parameters $\varphi$ and $\phi$ of the encoder and decoder, respectively, is formulated as the Evidence Lower Bound (ELBO)[34]. This takes the form
$\mathcal{L}_{VAE}(\mathbf{x},\varphi,\phi)=-\mathbb{E}_{q_{\phi}(\mathbf{z}|\mathbf{x})}\left[\log p_{\varphi}(\mathbf{x}|\mathbf{z})\right]+\beta_{1}D_{KL}(q_{\phi}(\mathbf{z}|\mathbf{x})||p(\mathbf{z})),$ (1)
where the first term expresses the reconstruction error, and a second term acts as regularization, penalising differences between the prior distribution and the approximate posterior of the latent variables. The hyperparameter $\beta_{1}$ determines a possible extra prior regularization weight, allowing further computational flexibility during training. The expectation in the first term is not explicitly calculated, but a Monte Carlo approximation is taken with a single sample from the variational posterior. With this approximation, the reconstruction error takes either the form of cross-entropy or mean-squared error between an input and its reconstruction from the single sample for the Bernoulli and normal likelihoods, respectively.
Explaining-away VAE
The EA-VAE incorporates into the standard VAE an additional scalar latent variable $s\in\mathbb{R}$ , which is inferred simultaneously with the other latents, $\mathbf{z}$, in the encoder, but acts multiplicatively in the decoder. That is, the mapping $\mathbf{f}(\mathbf{z})$ in the generative model of a usual VAE is supplemented by $s$ such that the mean of the likelihood in the EA-VAE is obtained as $s\,\mathbf{f}(\mathbf{z})$ (Figure 1e, right).
Consequently, the extension of the latent space yields a recognition model jointly for $\mathbf{z}$ and $s$: $q_{\phi}(\mathbf{z},s|\mathbf{x})$. To easily accommodate different potential probability distribution classes for the variational posterior, this joint variational posterior was broken down into two independent variational posterior components, $q_{\phi}(\mathbf{z}|\mathbf{x})$ and $q_{\psi}(s|\mathbf{x})$. As the scaling variable $s$ is meant to account for joint modulation of intensities across pixels, a positive value is justified. Consequently, a reasonable choice for the prior for $s$ is a Gamma distribution. In such a case the variational posterior was also taken as a Gamma. We note however that we explored several alternatives, as discussed in following sections showing no qualitative difference in the results.
Since $\mathbf{z}$ and $s$ are independent random variables, we write the objective function for the EA-VAE as
$\mathcal{L}_{EA-VAE}(\mathbf{x},\varphi,\phi)=-\mathbb{E}_{q_{\phi}(\mathbf{z}|\mathbf{x})}\left[\log p_{\varphi}(\mathbf{x}|\mathbf{z},s)\right]+\beta_{1}D_{KL}(q_{\phi}(\mathbf{z}|\mathbf{x})||p(\mathbf{z}))+\beta_{2}D_{KL}(q_{\psi}(s|\mathbf{x})||p(s)).$ (2)
Importantly, in the case of mapping $s$ by a softplus or an exponential function, the invertible transformation ensures that the Kullback-Leibler (KL) divergence remains unchanged [35]. For reconstruction, both $\mathbf{z}$ and $s$ are sampled from their respective posteriors before decoding.
VAE for inference on natural images
In this work we focus on inference over gray-scale images. That is, the input $\mathbf{x}$ corresponds here to a vector of intensity values for each of the pixels in an image. We first implemented and trained a Sparse VAE [19] identical to the one described in [20]. To allow for flexible learning, the model is constructed using fully connected layers (Figure S1a). In order to foster sparsity, this model employs a multivariate uncorrelated Laplace distribution for the prior distribution with zero mean and unit scale. To ensure that the KL divergence in the ELBO is analytically tractable, we chose to use the Laplace distribution for the variational posterior too: $q_{\phi}(\mathbf{z}|\mathbf{x})=\textrm{Laplace}(\mathbf{z};\boldsymbol{\mu}^{(\mathbf{x})},\boldsymbol{b}^{(\mathbf{x})})$, with $\boldsymbol{\mu}^{(\mathbf{x})}$ and $\boldsymbol{b}^{(\mathbf{x})}$ being the respective mean and scale parameters. These vectors are computed through amortized inference using the encoder for each specific image $\mathbf{x}$. We apply a standard sum of squared errors reconstruction loss, which corresponds to a normal observation noise assumption. That is, the likelihood for a given $\mathbf{x}$ is parameterized by a multivariate uncorrelated Normal distribution $p_{\varphi}(\mathbf{x}|\mathbf{z})=\mathcal{N}(\mathbf{x};f(\mathbf{x}),\boldsymbol{\sigma}_{obs}^{2})$, with $f(\mathbf{x})$ the mean. Thus, sampling the variational posterior, we obtain a reconstruction $\hat{\mathbf{x}}$ and Equation 1 takes the form of
$\mathcal{L}_{VAE}(\mathbf{x},\varphi,\phi)=$ $\frac{1}{2\sigma_{obs}^{2}}\sum_{i=1}^{M}(\hat{\mathrm{x}}_{i}-\mathrm{x}_{i})^{2}+\frac{M}{2}\log{(2\pi\sigma_{obs}^{2})}$ (3)
$+\beta_{1}\sum_{j=1}^{D}\left(-1+|\mu^{(\mathbf{x})}_{j}|-\log{(b_{j}^{(\mathbf{x})})}+b_{j}^{(\mathbf{x})}\exp{(-\frac{|\mu^{(\mathbf{x})}_{j}|}{b^{(\mathbf{x})}_{j}})}\right).$
It is worth mentioning that in cases when the observation noise is assumed to be drawn from an uncorrelated Normal distribution, then Equation 1 can be rewritten by global multiplication of $2\sigma_{obs}^{2}$ so that the $\beta$ hyperparameters absorbe the assumed observation noise [36], that is
$\beta^{\prime}=2\sigma_{obs}^{2}\beta.$ (4)
Further normalization can as well turn the first term in Equation 1 into the mean squared error (MSE). Selecting a larger $\beta$ as done in $\beta$-VAEs then corresponds to performing inference under an enlarged observation noise, fostering disentanglement by use of an uncorrelated prior.
EA-VAE for inference on natural images
The EA-VAE preserves the previous architecture but incorporates a scalar multiplicative latent variable $s$. For natural images patches, this variable is drawn from any specific selected distribution and represents in our model the contrast of the patch, as in previous work modeling natural image patches employing GSMs[11]. We examined different solutions for the implementation of this new architecture, ranging from a homogeneous latent space to having two separate encoders for $\mathbf{z}$ and $s$ (Figure S1), which variants we expand on below.
Scaling variable modeled as a softplus activated Laplace distribution
The central model presented in the main text aims to include a scaling variable with the smallest possible update to the standard VAE architecture. Here $s$ is one of the regular latents, with the same family of probability distributions for both the prior and the variational posterior, but later transformed by a monotonous, invertible function to the positive half-space, by use of a Softplus function (Figure S1a). We refer to this version of the model as the homogeneous version. The loss function in Equation 2 takes the following form
$\mathcal{L}_{EA-VAE}(\mathbf{x},\varphi,\phi)=$ $\frac{1}{2\sigma_{obs}^{2}}\sum_{i=1}^{M}(\hat{\mathrm{x}}_{i}-\mathrm{x}_{i})^{2}+\frac{M}{2}\log{(2\pi\sigma_{obs}^{2})}$ (5)
$+\beta_{1}\sum_{j=1}^{D-1}\left(-1+|\mu^{(\mathbf{x})}_{j}|-\log{(b_{j}^{(\mathbf{x})})}+b_{j}^{(\mathbf{x})}\exp{(-\frac{|\mu^{(\mathbf{x})}_{j}|}{b^{(\mathbf{x})}_{j}})}\right)$
$+\beta_{2}\left(-1+|\mu^{(\mathbf{x})}_{s}|-\log{(b_{s}^{(\mathbf{x})})}+b_{s}^{(\mathbf{x})}\exp{(-\frac{|\mu^{(\mathbf{x})}_{s}|}{b^{(\mathbf{x})}_{s}})}\right),$
where $\mu_{s}$ and $b_{s}$ are the inferred parameters of the scaling variable posterior.
We note that the resulting distribution (Laplace transformed by Softplus activation) does not have tractable moments. Thus we approximated the true posterior mean with samples in Figure 1h, given that the inferred variance of $s$ was low except below the observation noise, where slightly different samples become similarly close to 0 due to softplus activation (Figure S3). Also note the substitution of $D$ in in the first term of Equation 3 for $D-1$ in Equation 5, since we keep the overall number of latent variables unchanged for a fair comparison.
Beyond the homogeneous implementation of the model, we also explored other implementations, as described below. We note that no qualitative difference was observed in the results, highlighting the robustness of the proposal.
Scaling variable modeled as a Log-normal distribution
[…]
Scaling variable modeled as a Gamma distribution
A Gamma distribution can be parametrized in terms of the shape and scale parameters $k$ and $\theta$ respectively. For simplicity, we assumed that posteriors $q_{\psi}(s|\mathbf{x})=\textrm{Gamma}(s;k,\theta^{(\mathbf{x})})$ have a fixed shape parameter $k=2$. Since $k$ is fixed, the posterior’s mean, mode and variance depend only on the inferred scale parameter $\theta$. For the prior distribution, the shape parameter was also fixed at $k=2$ and the scale parameter $\theta=1/\sqrt{2}$ was chosen such that the prior had unit variance $k\theta^{2}=1$. A constant shape parameter $k$ means that the posterior standard deviation is always proportional to the posterior mean. This choice is in line with Weber’s Law of perception [37], where the estimation error of a quantity is a constant fraction of its mean. Architecture details in Figure S1b and Table S1. Under these assumptions, Equation 2 can be expressed similar to Equation 5, with the second KL term based upon Gamma distributions instead of Laplace [38]. The results from this model were analogous to the model presented in the main text (Figure S3).
VAEs and EA-VAEs for classic computer vision domains
In contrast to the natural image models, we explored using a convolutional neural network (CNN) architecture to train models on MNIST, fashionMNIST, and chestMNIST images, since convolutional architectures are more commonly employed in these domains (Figure S1c). These consist of encoders and decoders built of both fully-connected and convolutional layers (Figure S1c). Further, as is standard in the ML field, the variational posterior employed was a multivariate uncorrelated Normal distribution $q_{\phi}(\mathbf{z}|\mathbf{x})=\textrm{Normal}(\mathbf{z};\boldsymbol{\mu}^{(\mathbf{x})},\boldsymbol{\sigma}^{2(\mathbf{x})})$, with $\boldsymbol{\mu}^{(\mathbf{x})}$ and $\boldsymbol{\sigma}^{2(\mathbf{x})}$ the respective mean and variance vectors. In the case of the EA-VAE the variational posterior of the multiplicative variable, $s$, was log-normal: $q_{\psi}(s|\mathbf{x})=\textrm{Log-normal}(s;\nu^{(\mathbf{x})},\xi^{2(\mathbf{x})})$. Priors $p(\mathbf{z})$ and $p(s)$ come from the same family as their respective variational posteriors but with zero mean and diagonal unit variance.
Since the Gaussian pixel noise assumption strongly deviates from actual pixel statistics in the case of MNIST (which consists of black and white pixels), we assumed a bernoulli distribution in the likelihood. Accordingly, the cross-entropy reconstruction loss function was employed. In the case of cMNIST, cFashionMNIST and cChestMNIST contrast augmented datasets, the likelihood was parameterized by a multivariate uncorrelated Normal distribution, therefore resulting in mean squared error reconstruction error. The cost function constructed for the MNIST and ChestMNIST models were
$\mathcal{L}_{VAE}(\mathbf{x},\varphi,\phi)=\sum_{i=1}^{M}-(\mathrm{x}_{i}\log(\hat{\mathrm{x}}_{i})+(1-\mathrm{x}_{i})\log(1-\hat{\mathrm{x}}_{i}))+\beta^{\prime}_{1}\sum_{j=1}^{D}\frac{1}{2}\left(-1+(\mu^{(\mathbf{x})}_{j})^{2}-\log{(\sigma_{j}^{2(\mathbf{x})})}+\sigma_{j}^{2(\mathbf{x})}\right),$ (6)
$\mathcal{L}_{EA-VAE}(\mathbf{x},\varphi,\phi)=$ $\sum_{i=1}^{M}-(\mathrm{x}_{i}\log(\hat{\mathrm{x}}_{i})+(1-\mathrm{x}_{i})\log(1-\hat{\mathrm{x}}_{i}))+\beta^{\prime}_{1}\sum_{j=1}^{D-1}\frac{1}{2}\left(-1+(\mu^{(\mathbf{x})}_{j})^{2}-\log{(\sigma_{j}^{2(\mathbf{x})})}+\sigma_{j}^{2(\mathbf{x})}\right)$ (7)
$+\beta^{\prime}_{2}\frac{1}{2}\left(-1+(\nu^{(\mathbf{x})}_{j})^{2}-\log{(\xi_{j}^{2(\mathbf{x})})}+\xi_{j}^{2(\mathbf{x})}\right),$
while for the cMNIST, cFashionMNIST and cChestMNIST, the first term was correspondingly replaced by the sum of squared errors loss, that is $\sum_{i=1}^{M}(\hat{\mathrm{x}}_{i}-\mathrm{x}_{i})^{2}$.
Datasets
Natural Images
[…]
MNIST and ChestMNIST datasets
VAEs and EA-VAEs were trained on each domain separately. Each original $28\times 28$ pixelsize image was resized to $32\times 32$ due to the 3 layered convolutional encoder and decoder architecture, and the pixel intensity distribution. The pixel intensities for these canonical datasets are in the range $[0,1]$.
Contrast augmented cMNIST, cChestMNIST and cFashionMNIST datasets
Augmented versions of classic MNIST, ChestMNIST and FashionMNIST datasets were synthetically generated in order to have a rich variety in image contrast. Each image in the original dataset was first subtracted by $0.5$, so the pixel intensities move to the range $[-0.5,0.5]$. Then multiplied by a constant $c$ sampled from a Log-normal distribution, and then intensity was corrupted with a white noise vector $\boldsymbol{\eta}$, sampled from a Gaussian distribution of null mean and $\sigma_{obs}$ standard deviation
$\mathbf{x}^{aug}=c\left(\mathbf{x}^{ori}-\frac{1}{2}\right)+\boldsymbol{\eta}.$ (8)
This process was repeated such that cMNIST, cChestMNIST and cFashionMNIST datasets are 10 times larger than their original forms. $\sigma_{obs}$ values for each dataset are detailed in Table S2.
Shuffled pixels datasets
For testing purposes in the out of distribution experiments, all testing subsets (Van Hateren, MNIST, ChestMNIST, cMNIST, cChestMNIST, cFashionMNIST) were used to generate synthetically pixel shuffled datasets. Given a specific testset of size $(N_{img},N_{pix})$, the pixel-shuffled dataset is generated by performing random permutation of pixels such that pixels will get permuted only with pixels that are in the same location. By doing this, the pixel-shuffled dataset would have the same likelihood of observation, yet the information present in the shuffled image will differ from the original datasets, constituting a useful baseline for out-of-distribution detection.
Models’ training details
All models were trained in Python with Pytorch, either on a GPU NVIDIA TITAN V with a 64GB RAM memory (PC: 1) or on a NVIDIA GeForce RTX 2080 Ti with a 64GB RAM memory (PC: 2). Each model’s hyperparameters are detailed in Table S1 and Table S2.
For the standard VAE, we achieved the best results with beta-annealing: we set $\beta_{1}=0.01$ for the first 100 epochs, linearly increasing to the final $\beta_{1}=1$ over the next 100 epochs. Regarding Softplus-Laplace and Lognormal-Laplace EA-VAE models on natural images, $\beta_{1}=1$ was fixed during the whole training. To avoid collapse in the initial phase, a stronger regularization was applied on the scaling variable in these models for the course of the first 100 epochs ($\beta_{2}=10$), which linearly decreased to $\beta_{2}=1$ during the next 100 epochs. $\beta_{1}$ hyperparameter for the z-scored Gamma-Laplace model was fixed exploratorily such that the decoder’s reconstructions did not present visibly strong deviations from the original input. Results are robust with respect to changes in these hyperparameters, with no abrupt qualitative changes. For cMNIST, cChestMNIST and cFashionMNIST models, $\beta_{1}$ and $\beta_{2}$ were fixed by Equation 4 given the observation noise in the augmentation process.
All models were trained with early stopping so that the parameters were finally set at the epoch with best performance on the validation images. As a reference, the time taken to achieve each respective best epoch on PC 1 in this setting was of $62.2$ hours for the VAE vs $67$ hours for the EA-VAE trained on Van Hateren with Gamma distribution scaling variable. That is, the improvement in uncertainty representation in the EA-VAE comes at an expense of roughly an additional 10% computational time.
In the case of models trained on MNIST and ChestMNIST, although the EA-VAE has $23\%$ more parameters than the VAE, all models took around one hour to complete the 500 epochs. We note that the use of convolutions in this case drastically reduces training times from the case of natural images employing all fully connected layers.
Evaluation of VAEs and EA-VAEs trained on van Hateren
Latent receptive fields
The receptive field $\mathbf{RF}_{j}$ associated to the $j$-th latent dimension was computed through the Spike Triggered Average (STA) technique on the test dataset for all $j=1,...,1800$. The STA was estimated as
$\mathbf{RF}_{j}=\underset{\left\{\mathbf{x}\right\}}{\textrm{mean}}\,\mu_{j}^{(\mathbf{x})}\mathbf{x},$
where $\mu_{j}^{(\mathbf{x})}$ is the $j$-th component of the mean of $\mathbf{x}$’s inferred posterior $q_{\phi}(\mathbf{z}|\mathbf{x})$.
Latent units were discerned as informative or non-informative according to their contribution to the dataset reconstruction. When computing the relative change in the images reconstructed when all units are active, or when a specific unit is replaced by its mean activation throughout the ensemble, it is seen that the units associated to orientation-sensitive filters are clustered with the most contribution, while the second group barely contribute. All results shown in the main text were computed using only the informative latent dimensions.
Uncertainty representation in the latent space
We quantified the response of neurons for different observations and levels of uncertainty through the signal mean, signal variance and noise variance as a function of contrast.
For this, given an ensemble of patches $\left\{\mathbf{x}\right\}$, we infer the contrast $s^{*(\mathbf{x})}$ for each patch as the across-pixel standard deviation of intensities: $s^{*(\mathbf{x})}=\sigma_{pix}^{(\mathbf{x})}$.
The signal mean is defined as the average distance of the posterior’s mean to the prior’s mean (0 in this case) over all patches that have the same inferred contrast,
$\text{SM}(s)=\underset{\left\{\mathbf{x}|s^{*(\mathbf{x})}=s\right\}}{\textrm{mean}}\left\|\boldsymbol{\mu}^{(\mathbf{x})}\right\|.$
Signal variance is computed as the variance of the posterior’s mean of patches of same contrast, averaged across latent dimensions,
[…]
Finally, noise variance is defined as the posterior’s variance vector $\boldsymbol{\sigma}^{2(\mathbf{x})}$ averaged across latent dimensions and observations of equal inferred contrast $s$,
[…]
These quantities were computed for the test set, where no two patches have the same exact contrast. Hence, these were binarized with small enough bins such that no jumps were visible in the plots, and with error bars comparable to line-widths. The binarized expressions are
$\text{SM}(s_{i})=\underset{\left\{\mathbf{x}|s^{*(\mathbf{x})}\in B_{i}\right\}}{\textrm{mean}}\left\|\boldsymbol{\mu}^{(\mathbf{x})}\right\|$
[…]
The associated errors can be expressed as
$\epsilon_{{}_{\text{SM}}}(s_{i})=\frac{1}{\sqrt{N_{i}}}\sqrt{\underset{\left\{\mathbf{x}|s^{*(\mathbf{x})}\in B_{i}\right\}}{\textrm{Var}}\left\|\boldsymbol{\mu}^{(\mathbf{x})}\right\|}$
[…]
$\epsilon_{{}_{\text{NV}}}(s_{i})=\frac{1}{\sqrt{N_{i}}\sqrt{D}}\sqrt{\underset{\left\{j\cup\mathbf{x}|s^{*(\mathbf{x})}\in B_{i}\right\}}{\textrm{Var}}\boldsymbol{\sigma}^{2(\mathbf{x})}}.$
It is worth mentioning that the natural scenes dataset used in this work has patches with varied values of pixel intensity standard deviation $\sigma_{pix}^{(\mathbf{x})}$. As patches with high contrast live far away from the origin in the in the N-dimensional pixel space (since $s_{pix}^{(\mathbf{x})}=\sqrt{\frac{1}{N-1}\sum_{i=1}^{N}(x_{i}-\bar{x})^{2}}=\sqrt{\frac{1}{N-1}}\left\|\mathbf{x}\right\|$), it is expected that the network will have more difficulty exploring farther regions in the pixel space than the regions where low-contrast patches live. Therefore, in evaluation instances of the Gamma-Laplace model, patches with contrast higher than $\sigma_{pix}=2$ were not considered.
Divisive normalization experiments
To characterize the dependencies between latents of the Variational Autoencoders, we calculated conditional distributions for pairs of latents by analyzing the distribution of posterior means for a large number of natural image patches. Across a large number of randomly sampled filter pairs the variance of the conditional distribution was calculated at four quadrants: two central and two flanking quadrants (Figure 5a), which corresponded to low-level, and high-level activations of the conditioned filter, respectively (Figure 5b).
We modeled divisive normalization by predicting responses of individual latents of the Variational Autoencoder through the linear filter response of that latent divided by the weighted sum of the (linear) responses of all other latents
$z_{i}=\frac{L_{i}}{\sqrt{\sum_{j\neq i}w_{ji}L_{j}^{2}+\sigma^{2}}}.$ (9)
Here $L_{j}$ is the linear response of latent variable $j$ and weights are constrained to be positive. Weights, $w_{ji}$ were determined through regression. Linear responses of latents were calculated by taking the dot product of the image and the vector mapping the latent to the pixel space in the generative model. Fitting of the model was driven by the deviation of the linear responses from the posterior produced by the recognition model.
Evaluation of VAEs and EA-VAEs trained on MNIST and ChestMNIST
Uncertainty of latent representations
For all experiments involving uncertainty in the latent space, we compute the uncertainty of the network for a given input $\mathbf{x}$ as the standard deviations of the posterior $q_{\phi}(\mathbf{z}|\mathbf{x})$ averaged across latent dimensions,
[…]
We call this as well the posterior width.
Uninformative observations
We evaluated the models’ ability to handle uninformative observations by creating an "average image" through pixel-wise averaging of all dataset images. This image served as a proxy for an uninformative observation. For each model (VAE and EA-VAE), we characterized latent uncertainty by Equation 10.
Morphing and uncertainty
We conducted an experiment to study the uncertainty in the VAE and EA-VAE latent space when morphing digits from one label to another. The first step for doing this was selecting all the test images that belong to a specific label A and computing the average coordinate in the latent space of these (Figure S4a bottom panels). The output $\hat{\mathbf{x}}_{A}$ of the decoder from that position is a representative image of the label A (Figure S4a top). The same procedure is repeated with images belonging to label B obtaining $\hat{\mathbf{x}}_{B}$. Then, morphing is done by linearly combining $\hat{\mathbf{x}}_{A}$ and $\hat{\mathbf{x}}_{B}$ with weight $\lambda$ varying from 0 to 1, that is
$\hat{\mathbf{x}}_{AB,\lambda}=(1-\lambda)\hat{\mathbf{x}}_{A}+\lambda\hat{\mathbf{x}}_{B}.$
We then inferred the corresponding posteriors taking these morphed images as input, and computed the uncertainty of the network by Equation 10,
[…]
Through this procedure, uncertainty curves were computed for all the combinations of morphing from labels 00 to $9$ (Figure S4b color curves).
To evaluate if a certain curve has the expected behavior of being maximally uncertain at intermediate $\lambda$ values, we fitted a second order polynomial function (Figure S4c gray curves) and checked whether it’s curvature was negative and the maximum was located inside a window of size $L$ centered respect to $\lambda=0.5$, for different $L$ values (Figure S4c).
Image corruption
The image corruption experiment was done over the complete test set. Each image was synthetically corrupted by applying a two-dimensional Gaussian kernel of covariance $\eta_{b}I_{2x2}$. The resulting blurred image was encoded into the models’ latent space, where the measure of uncertainty was computed. We compute the uncertainty of the network for a given input by Equation 10.
This process was repeated for all images, and the average uncertainty was computed. This was repeated for different $\eta_{b}$ values until the images were completely blurred out ($\eta_{b}=16$ equal to half the pixel width of the images) We repeated the experiment with pixel-noise corruption. In this case, each pixel was intensity corrupted by additive white Gaussian noise of null mean and variance $\eta_{p}^{2}$. The whole perturbed image was then rescaled in intensity to preserve the original range of intensities.$\eta_{p}$ values between 0 and 1 were considered, going from no perturbation to a strongly perturbed image.
Out of distribution experiments
To study the capacity of the trained models to associate near-to-prior uncertainty for images from previously unseen distributions, we compared the latent representations of images in and out of the distribution used for training. The reconstruction quality was analysed by projecting the posterior mean back to the pixel-space using the decoder to observe how OOD inputs were represented. Latent uncertainty to a given observation was computed the same way as in the two previous experiments. We looked at models trained with all datasets (MNIST, ChestMNIST, cMNIST, cFashionMNIST, cChestMNIST and Van Hateren) and tested with the complementary domains and corresponding pixel-shuffled dataset (Figure 2c and Figure S5). In all cases, images from the OOD datasets were resized to coincide the size of images from the training set, and rescaled in pixel intensity to match the in-distribution pixel intensity distribution. For the models trained on Van Hateren, that is population wise z-scoring the OOD datasets. For all other models trained on MNIST domains, that is rescaling the Van Hateren’s dataset pixel intensity distribution such that it’s mean $\pm$ 3 deviations fit the range [0,1].
We found the same behavior as shown in the main text when training models with x-ray images and testing on natural images (Figure S5d, rightmost block). Yet when testing on digits and fashion images, both models seem to fare comparably, with reported uncertainties that do not approach the prior, although the reconstructed images do not collapse to within domain examples. When training with natural images and testing on the other (Figure S5e), the distribution of reported uncertainties, which originally had two modes for within distribution images corresponding to more informative (high contrast) and less informative (low contrast) images, reduces to a single mode mode for VAEs and EA-VAEs. The asymmetry in the results when original and target domains are inverted suggests a hierarchy in the richness of the feature space learned from the statistics of the images.
Predictive uncertainty
To evaluate the role of uncertainty representations in classification, two separate MLPs were trained to classify MNIST handwritten digits into labels using samples from the posterior latent representations of VAEs and EA-VAEs. Once trained, the classification task involved two experiments. In the first, a set of continuously morphed digits was used to analyze predictive uncertainty, measured as the entropy of output distributions. We compared the entropy curves of the MLPs’ predictions by summing the entropy for all intermediate images for the VAE versus the EA-VAE. A paired t-test quantified the differences. In the second experiment, we tested the MLPs with out-of-distribution inputs, including ChestMNIST and pixel-shuffled MNIST images, and compared the entropy of predictions based on the two latent representations.
MNIST MLP classifier
A multilayer perceptron (MLP) was trained to classify MNIST digits into its labels. Given a digit $\mathbf{x}$, the network receives as an input a sample $\mathbf{z}$ of the posterior distribution $q_{\phi}(\mathbf{z}|\mathbf{x})$ encoded by the already trained and fixed VAE (or EA-VAE), and outputs a 10-dimensional vector quantifying the predicted probability for each label $l$: $\mathbf{w}=\left[P(l=0),P(l=1),...,P(l=9)\right]$. The MLP had two hidden layers of 50 units with relu non-linearity. For the output layer, softmax activation function was used. The model was trained for 1000 epochs, where for each epoch, resampling of the posterior distributions was done for computing the input of the network. Adam optimizer was used.
Data availability
All datasets used in this paper are publicly available. Van Hateren dataset of natural image patches of 40x40 pixels is provided in Ref. [20]. MNIST and FashionMNIST datasets can be downloaded via Pytorch, and ChestMNIST dataset can be downloaded and imported using the MEDMNIST official repository.
Code availability
Implementation code and instructions for reproducing results are freely available on GitHub [40].
Acknowledgments
This work was supported by Argentina’s National Scientific and Technical Research Council (CONICET) (RE), which covered the salaries of all the authors, the the European Union project RRF-2.3.1-21-2022-00004 within the framework of the Artificial Intelligence National Laboratory (GO) and the National Research, Development and Innovation Office under grant agreement ADVANCED 150361 (GO). The authors gratefully acknowledge NVIDIA Corporation for the GPU computing, and the support of Universidad Nacional del Litoral (Grants CAID-PIC-50220140100084LI, 50620190100145LI), ANPCyT (PICT-PRH-2019-00009, 2022-03-00593).
Supplementary Information
Fig. S1: a, Trained VAE (left) and EA-VAE (right) models for inference on natural images. The encoder’s input of size $M=1600$ is followed by two fully connected hidden layers with softplus activation function. Two separate fully connected layers of $D=1800$ units encode the mean and variance of $q_{\phi}(\mathbf{z}|\mathbf{x})$. For the VAE, one fully connected layer decodes the latent sample into a reconstructed patch. For the EA-VAE model, the described VAE was modified such that the last component of the sample of the posterior distribution passes through function g() defining the effective $s$ sample from $q_{\psi}(s|\mathbf{x})$. The posterior sampled $s$ value is globally multiplied with the decoder’s output, resulting in the final reconstructed patch. b, Trained VAE (left) and EA-VAE (right) models for inference on natural images when scaling variable assumed from a Gamma distribution. In this case, the scaling variable is inferred from a separate encoder. $k=2$ was chosen as the smallest integer number that qualitatively achieved a similar distribution to the dataset’s $\sigma_{pix}$ distribution. Computationally, the choice of a fixed integer $k$, simplifies the sampling procedure, since one can always sample from the same standard Gamma function and then reparameterize by the scale parameter. This enables simple gradient computations across the sampling process. c, Trained VAE (left) and EA-VAE (right) models for inference on MNIST classic and contrast-augmented domains (digits, chest x-rays and fashion). For the VAE, the encoder’s input of size $M=32$ is followed by three convolutional hidden layers with relu activation function. Two separate fully connected layers of $D$ units encode the mean and variance of $q_{\phi}(\mathbf{z}|\mathbf{x})$. A fully connected and 3 deconvolutional layers decode the latent sample. Output from last layer passes trough h() function giving place to the reconstructed image. For the EA-VAE, the described VAE model in was modified such that separate fully connected layers encode mean and variance of a Normal distribution. Its sample passes through function g() defining the effective $s$ sample from $q_{\psi}(s|\mathbf{x})$. The posterior sampled $s$ value is globally multiplied with the decoder’s output before going through function g(), resulting in the final reconstructed image.
Van Hateren Model PC $D$ g( ) $\sigma_{obs}$ $\beta^{\text{VAE}}_{1}$ $\beta_{1}^{\text{EA-VAE}}$ $\beta_{2}^{\text{EA-VAE}}$ lr wd Epochs Batch Size
SoftLaplace-Laplace 2 1800 softplus 0.4 0.01 $\rightarrow$ 1 1 10 $\rightarrow$ 1 3e-5 1e-6 5000 512
LogNormal-Laplace 2 1800 exp 0.4 0.01 $\rightarrow$ 1 1 10 $\rightarrow$ 1 3e-5 1e-6 5000 512
Gamma-Laplace 1 1800 0.46/$\alpha$ 1 1 2 3e-5 1e-5 10000 128
($\alpha\approx 5.315$) $\approx 0.09$
Table S1: Hyperparameter settings for the Van Hateren models
Model PC $D$ h( ) g( ) $\sigma_{obs}$ $\beta^{\prime}_{1}$ $\beta^{\prime}_{2}$ lr Epochs Batch Size
MNIST 1 5 sigmoid exp 4 1 1e-3 500 128
ChestMNIST 1 5 sigmoid exp 1 1 1e-3 500 128
cMNIST10D 1 10 exp 0.25 0.125 0.125 1e-4 500 128
cMNIST3D 1 3 exp 0.25 0.125 0.125 1e-4 500 128
cFashionMNIST 1 10 exp 0.25 0.125 0.125 1e-4 500 128
cChestMNIST 1 4 exp 0.1 0.02 0.02 1e-4 500 128
Table S2: Hyperparameter settings for the MNIST domain models
Fig. S2: a, Example test image patches of increasing contrast levels, and their respective reconstruction through the Standard VAE and the SoftLaplace-Laplace EA-VAE. b, Extended examples of informative latent receptive filters in the standard VAE and the SoftLaplace-Laplace EA-VAE.
Fig. S3: Properties of inferred posteriors as contrast is varied for EA-VAE models trained on natural image patches for variations on how scaling variable is modeled: either from a LogNormal, Gamma or Softplus-activated Laplace distribution. a, Example test image patches and their respective reconstruction through EA-VAE. b, Receptive filters of example latents in the EA-VAEs. c, Signal mean, signal variance and noise variance of the inferred latent posteriors in natural image-trained EA-VAEs as a function of image contrast. Prior mean and variance are shown in horizontal dashed black lines. Observation noise assumed in the model is shown in vertical dotted black lines. d, Inferred posterior mean and posterior std. of the scaling variable for individual patches (dots) in the EA-VAE model, as a function of the measured contrast of these images. Observation noise assumed in the model is shown in vertical dotted black lines.
Fig. S4: Morphing experiments in MNIST. a, Centers for the distribution corresponding to each digit are first found. b, Posterior widths for the morphed digits with standard VAE (left) and EA-VAE (right). Position where uncertainty is maximal is computed via quadratic fit. c, For a given window size around the middle point between the centers, the fraction of interpolations where the maximal uncertainty lies within this window is computed via b fittings.
Fig. S5: Out of distribution experiments. Same as in Fig. 2d, now when training on a, cMNIST 10D, b, cFashionMNIST 10D, c, cChestMNIST, d, ChestMNIST and e, Gamma-Laplace Van Hateren ($z$-scored).
Fig. S6: Behaviour of trained MLP classifier when gradually morphing between image samples from digit ’A’ to digit ’B’. Entropy of predicted probability distribution as a function of combination weight of labels for all 45 possible combinations of labels. Dashed line corresponds to the upper limit determined by the discrete uniform distribution.